How many gallons of a 10% salt solution must be mixed with 12 gallons of a 15% salt solution to obtain a 12% salt solution?
Answer by mananth(16949)

-------------- percent ---------------- Amount
Salt solution I 15 ---------------- 12 gallons
salt solution II 10 ----------------x gallons
Total 12 ---------------- 12 + x gallons

15*12+10x=12(12 +x)
180+10x =144+12 x
10x-12x =-180+144
-2x =-36
/ -2
x = 18 of gallons salt solution II
18 of gallons Salt solution I
m.ananth@hotmail.ca